# CookAR: Affordance Augmentations in Wearable AR to Support Kitchen Tool Interactions for People with Low Vision

### v1

@@ Abstract.
Cooking is a central activity of daily living, supporting independence and both mental and physical health. However, prior work has highlighted key barriers for people with low vision (LV) to cook, particularly around safely interacting with cooking tools, such as sharp knives or hot pans. Drawing on recent advancements in computer vision (CV), we present CookAR, a head-mounted AR system with real-time object affordance augmentations to support safe and efficient interactions with kitchen tools. To design and implement CookAR, we manually collected and annotated the first egocentric dataset of kitchen tool affordances, fine-tuned an affordance segmentation model, and developed an AR system with a stereo camera to generate visual augmentations. To validate CookAR, we conducted a technical evaluation of our fine-tuned model as well as a qualitative lab study with 10 LV participants for suitable augmentation design. Our technical evaluation demonstrates that our model outperforms the baseline on our tool affordance dataset, while our user study indicates a preference for affordance augmentations over the traditional whole object augmentations. Code is available at: https://github.com/makeabilitylab/CookAR.
@@ 1. Introduction
Cooking is an essential activity of daily living, supporting independence (Miyawaki et al., 2012; Mynatt and Rogers, 2001; Taylor et al., 2016) and both mental and physical health (Mills et al., 2017; Bilyk et al., 2009; Jones et al., 2019; Taylor et al., 2016). However, cooking also involves significant visual tasks that can be challenging or dangerous for blind and low vision (BLV) people, especially when interacting with kitchen tools, such as sharp knives or hot pans (Bilyk et al., 2009; Jones et al., 2019; Li et al., 2021, 2024; Wang et al., 2023b).
@@
Unlike those who are completely blind, people with low vision (LV)—vision loss that cannot be corrected using glasses or contact lenses (Corn and Erin, 2010)—often rely on their residual vision in daily activities and use different low vision tools to enhance visual information (Szpiro et al., 2016a, b). With recent advancements in AI-powered augmented reality (AR) technology, researchers have explored new possibilities for supporting LV individuals by automatically recognizing their environment and providing appropriate visual augmentations. Prior AR research prototypes have been developed to assist with tasks such as visual search (Zhao et al., 2016), stair navigation (Zhao et al., 2019), and sports (Lee et al., 2023). However, none have been specifically designed to support meal preparation. Moreover, prior AR systems mainly focus on understanding effective augmentation designs (Zhao et al., 2019), often oversimplifying the computer vision (CV) recognition in their development, thus neglecting the effects of technological limitations (e.g., CV inaccuracies and system delays) on user experience.
@@
We introduce CookAR, a fully-functional wearable stereo AR prototype that recognizes and augments the affordances of cooking tools in real-time to support LV individuals with efficient and safe interactions in the kitchen. In contrast to prior research that augments objects as a whole (Zhao et al., 2016; Fox et al., 2023), we distinguish and augment the object affordance (i.e., component parts that afford interactions), such as the safe-to-handle “grabbable” areas and the dangerous-to-touch “hazardous” areas (Figure 1). To enable accurate affordance recognition, we constructed a custom egocentric image dataset for kitchen tool affordances by selecting and labeling images from the Epic Kitchens dataset (Damen et al., 2018) and fine-tuned an affordance segmentation model. We then leveraged stereo depth estimation using the ZED Mini111https://store.stereolabs.com/products/zed-mini stereo camera and an Oculus Quest 2222https://www.meta.com/quest/products/quest-2/ headset with video passthrough to precisely overlay affordance augmentations on the 3D environment in real-time.
@@
To evaluate CookAR, we conducted a technical evaluation of our fine-tuned model as well as a three-part qualitative lab study with 10 LV participants to evaluate their experiences with CookAR and solicit feedback about potential affordance augmentation designs. For the model assessment, we found that our fine-tuned affordance segmentation model (mAP of 46.3%) outperformed the base RTMDet (Lyu et al., 2022) model (mAP of 12.3%) in accurate tool affordance recognition and segmentation. For the three-part user study, LV participants were first asked to locate and pickup cooking tools across three conditions: (1) without CookAR (baseline), (2) with CookAR displaying whole object augmentations, and (3) with CookAR displaying affordance augmentations. They then completed a free-form cooking task with CookAR (Part 2) and brainstormed desired augmentation designs using design probes (Part 3). Findings indicate that participants prefer affordance augmentations over whole object augmentations in a kitchen, as they enable faster understanding of an object’s spatial arrangement and safe interaction parts. Most participants preferred affordance augmentations consisting of green solid overlay on grabbable areas and red outlines on hazardous areas. Moreover, participants identified five additional tool affordances with desired augmentations, including entry (e.g., cup rim), exit (e.g., carafe spout), containment (e.g., cup base), intersection (e.g., knife blade on butter), and activation (e.g., carafe buttons) areas, all of which should be outlined in a contrasting color (e.g., black or white).
@@
In summary, our research contributions include: (1) CookAR, a fully-functional AI-powered wearable AR prototype that augments kitchen tool affordances for low vision users to enable safe and efficient tool interactions; (2) an egocentric affordance dataset for kitchen tools and an accompanying fine-tuned affordance segmentation model; and (3) user study results with 10 LV participants that reveal user experiences with CookAR, preferences for augmentation designs, and five newly desired affordance areas.
@@ 2. Related Work
Our work builds on prior formative research exploring low-vision cooking, wearable AR to enhance accessibility, and affordance segmentation.
@@ 2.1. Challenges in Low Vision Cooking
Low vision people face various challenges in activities of daily living, such as cooking (Bilyk et al., 2009; Wang et al., 2023a, b; Li et al., 2021, 2024), shopping (Szpiro et al., 2016b), navigation (Zhao et al., 2018; Szpiro et al., 2016b), and sports (Lee et al., 2023; Rector et al., 2015). Among these activities, cooking is an essential task for an independent and healthy life (Taylor et al., 2016). However, cooking also poses major barriers and safety concerns to blind and low vision (BLV) people as they need to interact with various ingredients and kitchen tools, such as sharp knives and hot pans (Kashyap, 2020; Kim et al., 2022). As a result, BLV people tend to eat more pre-processed food or frequently dine at restaurants, which can negatively impact their health (Jones et al., 2019; Montero, 2005).
@@
To better understand how BLV people engage in cooking tasks, prior work has conducted both interview and observational studies (Li et al., 2021; Lin et al., 2014; Wang et al., 2023a, b). For example, Jones et al. (Jones et al., 2019) surveyed 101 BLV participants in the U.K. about their shopping and cooking experiences, revealing that vision loss made cooking difficult and the difficulty level was correlated with the severity of visual impairments. Li et al. (Li et al., 2021) analyzed 122 YouTube videos of BLV people preparing meals and interviewed 12 BLV participants about their cooking experiences. They identified several cooking-related challenges, such as utilizing cooking tools and tracking object dynamics in the kitchen. A follow-up contextual inquiry study (Li et al., 2024) examined how BLV people recognize cookware and utensils and measure ingredients. The study identified essential cooking-related information to convey, such as position, safety, and orientation information.
@@
Specifically for low-vision (LV) people, Wang et al. (Wang et al., 2023a) conducted a contextual inquiry study, observing and comparing the cooking experiences between six LV participants and four blind participants. They found that while blind participants relied on touch, LV people extensively used their vision during cooking. However, compared to blind people, LV people felt less confident, less safe, and more tired and stressed during cooking. Moreover, they were less satisfied with the currently-available cooking tools than blind people, indicating a need for technology that considers low vision people’s unique needs. The study also identified key challenges faced by LV people, such as distinguishing objects with low contrast and safely interacting with dangerous kitchen tools. Wang et al. (Wang et al., 2023b) further interviewed six low vision rehabilitation professionals to understand current training strategies and tools to support cooking. However, they highlighted that current solutions cannot fully address all cooking challenges faced by LV people. Our research fills this gap by building an AR system to support safe and efficient cooking tool interactions for low vision people by providing visual augmentations.
@@ 2.2. Using Wearable AR to Enhance Accessibility
In the fields of accessibility and HCI, wearable AR has been used to support people with diverse disabilities, for example, AR systems that caption and visualize speech and sounds for deaf or hard of hearing (DHH) individuals (Findlater et al., 2019; Jain et al., 2018a, b, 2015; Miller et al., 2017; Olwal et al., 2020; Peng et al., 2018; Schipper and Brinkman, 2017), support hands-free interactions with screen displays for people with upper body motor impairments (McNaney et al., 2014; Malu and Findlater, 2014, 2015), offer speech support for people with aphasia (Williams et al., 2015), and provide social cue therapy for children with autism (Washington et al., 2017).
@@
Within the low-vision aid context, head-worn AR devices can selectively enhance user’s vision by interpreting their environment and tasks (Azenkot and Zhao, 2017; Zhao et al., 2017). For instance, prior research has developed AR systems that can capture real-time video feeds of the surroundings and apply image processing techniques (Massof and Rickman, 1992; Zhao et al., 2016; Deemer et al., 2018) to enhance visual information, such as edge enhancement (Massof et al., 1994; Kwon et al., 2012; Hwang and Peli, 2014), scene recoloring based on distance (Everingham et al., 1999; Hicks et al., 2013; van Rheede et al., 2015), and pixel remapping for visual field loss (Loshin and Juday, 1989; Gupta et al., 2018; Peli, 2001; Luo and Peli, 2006). However, while these solutions are beneficial for simple tasks like reading (Stearns et al., 2017; Huang et al., 2019; Deemer et al., 2018), they still lack semantic understanding of the scene and cannot effectively support more complex activities involving object interactions, such as cooking (Wang et al., 2023b).
@@
More recently, researchers have combined AR and CV to develop scene-aware visual augmentations aimed at assisting LV people with more intricate activities like visual search (Zhao et al., 2016), stair navigation (Zhao et al., 2019), wayfinding (Zhao et al., 2020), obstacle maneuver (Fox et al., 2023), button pressing (Lang and Machulla, 2021), and sports (Lee et al., 2023). Nonetheless, no prior systems have addressed the unique cooking challenge that involves dynamic tool interactions. Moreover, prior AR research for low vision mainly focuses on designing and evaluating suitable visual augmentations for low vision people. They tend to oversimplify the CV recognition in system development, for example, using QR codes (Zhao et al., 2016, 2019) or existing spatial mapping APIs (Zhao et al., 2020; Fox et al., 2023) to anchor augmentations to the real-world environment. As a result, they overlook the technical challenges in building a real-time AI-powered AR system, as well as the potential impact of the technological limitations (e.g., CV errors, system latency) on users’ experiences. Our research addresses this gap by developing CookAR, a fully-functional wearable AR system that recognizes and augments kitchen tool affordances in real-time to support safe and efficient interactions in a cooking scenario.
@@ 2.3. Affordance Segmentation
In contrast to most prior research that augments whole objects (Zhao et al., 2016; Fox et al., 2023; Lee et al., 2023), our research focuses on recognizing and augmenting tool affordances. Affordance is traditionally defined as “the opportunities for actions that objects offer, relative to the user’s ability to perceive and act on them” (Gibson, 1966, 1977, 2014). Highlighting object affordances can effectively guide human attention and actions (Garrido-Vásquez and Schubö, 2014; Tünnermann et al., 2015; Rehrig et al., 2022). Despite the prominence of affordance segmentation in robotics (Chu et al., 2019a, b; Nguyen et al., 2016, 2017; Bahl et al., 2023) and computer vision (Luddecke and Worgotter, 2017; Chuang et al., 2018; Luo et al., 2022), automatic affordance augmentation has received comparatively less attention and applications in the field of HCI. Notably, there exists a gap for an affordance dataset in the context of a kitchen environment, especially designed for the needs of individuals with low vision. To address this, we built an AR system that focuses on affordance segmentation and enhancement. We created a new dataset focused on the affordances of kitchen tools by selecting and annotating image frames from the egocentric Epic Kitchens dataset (Damen et al., 2018) and fine-tuned an instance segmentation model on our dataset.
@@ 3. System Implementation
We designed and built CookAR, an innovative solution that integrates AR and CV technologies to improve object interactions for individuals with low vision. Unlike traditional enhancements that target objects as a whole, our prototype highlights their affordances (i.e., functional parts), facilitating easier identification and interaction with areas to grasp or avoid. To create a fully-functional wearable AR system, we needed to address both the computer vision problem of accurately recognizing object affordances in real-time and the HCI problem of designing and rendering suitable affordance augmentations. In this section, we describe our approach for each step, including (1) collecting and annotating a dataset focused on the affordances of kitchen tools; (2) fine-tuning an instance segmentation and recognition model on our dataset to detect these affordances; and (3) building a head-mounted AR system with a stereo camera to render visual augmentations based on the outputs of our affordance recognition model.

### v2

@@ Abstract.
Cooking is a central activity of daily living, supporting independence as well as mental and physical health. However, prior work has highlighted key barriers for people with low vision (LV) to cook, particularly around safely interacting with tools, such as sharp knives or hot pans. Drawing on recent advancements in computer vision (CV), we present CookAR, a head-mounted AR system with real-time object affordance augmentations to support safe and efficient interactions with kitchen tools. To design and implement CookAR, we collected and annotated the first egocentric dataset of kitchen tool affordances, fine-tuned an affordance segmentation model, and developed an AR system with a stereo camera to generate visual augmentations. To validate CookAR, we conducted a technical evaluation of our fine-tuned model as well as a qualitative lab study with 10 LV participants for suitable augmentation design. Our technical evaluation demonstrates that our model outperforms the baseline on our tool affordance dataset, while our user study indicates a preference for affordance augmentations over the traditional whole object augmentations.
@@ 1. Introduction
Cooking is an essential activity of daily living, supporting independence (Miyawaki et al., 2012; Mynatt and Rogers, 2001; Taylor et al., 2016) as well as both mental and physical health (Mills et al., 2017; Bilyk et al., 2009; Jones et al., 2019; Taylor et al., 2016). However, cooking also involves significant visual tasks that can be challenging or dangerous for blind and low vision (BLV) people, especially when interacting with kitchen tools, such as sharp knives or hot pans (Bilyk et al., 2009; Jones et al., 2019; Li et al., 2021, 2024; Wang et al., 2023b).
@@
Unlike those who are completely blind, people with low vision (LV)—vision loss that cannot be corrected using glasses or contact lenses (Corn and Erin, 2010)—often rely on their residual vision in daily activities and use different low vision tools to enhance visual information (Szpiro et al., 2016a, b). With recent advancements in AI-powered augmented reality (AR), researchers have explored new possibilities for supporting LV individuals by automatically recognizing their environment and providing appropriate visual augmentations, including stair navigation (Zhao et al., 2019), visual search (Zhao et al., 2016), and sports (Lee et al., 2023). While promising, these previous AR systems focus primarily on understanding effective augmentation designs (Zhao et al., 2019), often oversimplifying the computer vision (CV) recognition in their development, thus neglecting the effects of technological limitations (e.g., CV inaccuracies and system delays) on user experience. Moreover, though there has been substantial formative work in BLV cooking within the HCI literature (Li et al., 2021, 2024; Wang et al., 2023b), no previous AR system has been built to address LV cooking specifically.
@@
We introduce CookAR, a wearable stereo AR prototype that recognizes and augments cooking tool affordances in real-time to support LV meal preparation. In contrast to prior AR research that highlights objects as a whole (Zhao et al., 2016; Fox et al., 2023), we distinguish and augment the object affordance specifically (i.e., component parts that afford interactions), such as the safe-to-handle “grabbable” areas and the dangerous-to-touch “hazardous” areas (Figure 1). To enable accurate affordance recognition, we constructed a custom egocentric image dataset for kitchen tool affordances by selecting and labeling images from the Epic Kitchens dataset (Damen et al., 2018) and fine-tuned an affordance segmentation model. We then combined the ZED Mini111https://store.stereolabs.com/products/zed-mini stereo camera and an Oculus Quest 2222https://www.meta.com/quest/products/quest-2/ headset to achieve a video passthrough AR system with CV and stereo depth estimation capabilities to precisely overlay affordance augmentations on the 3D environment in near real-time.
@@
To evaluate CookAR, we conducted a technical evaluation of our fine-tuned model as well as a three-part qualitative lab study with 10 LV participants. For the model assessment, we found that our fine-tuned affordance segmentation model (mAP of 46.3%) outperformed the base RTMDet (Lyu et al., 2022) model (mAP of 12.3%) in tool affordance recognition and segmentation. For the three-part user study, LV participants were first asked to locate and pick up cooking tools across three conditions: (1) with their typical method in daily life such as wearing corrective lenses (i.e., real-world baseline); (2) with CookAR displaying whole object augmentations (i.e., augmentation baseline); and (3) with CookAR displaying affordance augmentations. They then completed a free-form cooking task with CookAR (Part 2) and brainstormed desired augmentation designs using design probes (Part 3). Findings indicate that participants prefer affordance augmentations over whole object augmentations in a kitchen as they enable faster understanding of an object’s spatial arrangement and safe interaction parts. Most participants preferred affordance augmentations consisting of green solid overlay on grabbable areas and red outlines on hazardous areas. Moreover, participants identified five additional tool affordances with desired augmentations, including entry (e.g., cup rim), exit (e.g., carafe spout), containment (e.g., cup base), intersection (e.g., knife blade on butter), and activation (e.g., carafe buttons) areas—all which should be outlined in a contrasting color (e.g., black or white).
@@
In summary, our research contributions include: (1) CookAR, a fully-functional AI-powered wearable AR prototype that augments kitchen tool affordances for low vision users to enable safe and efficient tool interactions; (2) an egocentric affordance dataset for kitchen tools and an accompanying fine-tuned affordance segmentation model. To enable others to build off our work, this dataset and model are open-sourced at: https://github.com/makeabilitylab/CookAR; and (3) user study results with 10 LV participants that reveal user experiences with CookAR, preferences for augmentation designs, and five newly desired affordance areas.
@@ 2. Related Work
Our work builds on prior formative studies on low-vision cooking, wearable AR for accessibility, and affordance segmentation.
@@ 2.1. Challenges in Low Vision Cooking
People with low vision (LV) face challenges in everyday activities, such as cooking (Bilyk et al., 2009; Wang et al., 2023a, b; Li et al., 2021, 2024), shopping (Szpiro et al., 2016b), navigation (Zhao et al., 2018; Szpiro et al., 2016b), and sports (Lee et al., 2023; Rector et al., 2015). Among these, cooking is an essential task for an independent and healthy life (Taylor et al., 2016). However, this task also poses major accessibility barriers and safety concerns to BLV individuals, including interacting with sharp knives and hot pans (Kashyap, 2020; Kim et al., 2022). Consequently, they tend to eat more pre-processed food or frequently dine at restaurants, which can negatively impact their health (Jones et al., 2019; Montero, 2005).
@@
To better understand how BLV people engage in cooking tasks, prior work has conducted both interview and observational studies (Li et al., 2021; Lin et al., 2014; Wang et al., 2023a, b). For example, Jones et al. (Jones et al., 2019) surveyed 101 BLV participants in the U.K. about their shopping and cooking experiences, revealing that vision loss made cooking difficult and that the level of difficulty was correlated to the severity of visual impairments. Li et al. (Li et al., 2021) analyzed 122 YouTube videos of BLV people preparing meals and interviewed 12 BLV participants about their cooking experiences. They identified several cooking-related challenges, such as utilizing cooking tools and tracking object dynamics in the kitchen. A follow-up contextual inquiry study (Li et al., 2024) examined how BLV people recognize cookware and utensils and measure ingredients. This study highlighted essential cooking-related information to convey, such as position, safety, and orientation of objects.
@@
Specifically for LV people, Wang et al. (Wang et al., 2023a) conducted a contextual inquiry study, observing and comparing the cooking experiences between six LV participants and four blind participants. They found that while blind participants relied on touch, LV people used their vision extensively while cooking. However, compared to blind people, LV people felt less confident, less safe, and more tired and stressed due to their reliance on impaired vision. Moreover, LV individuals were less satisfied with existing cooking tools compared to blind people, highlighting the need for technology that considers their unique needs. The same study also identified key challenges LV people face, such as distinguishing objects with low contrast and safely interacting with dangerous kitchen tools. Wang et al. (Wang et al., 2023b) further interviewed six LV rehabilitation professionals to understand current training strategies and tools for cooking. They emphasized that existing solutions are insufficient to overcome all cooking challenges LV people experience. Our research fills this gap by creating an AR system that supports LV people in safely and efficiently using cooking tools through visual augmentations.
@@ 2.2. Using Wearable AR to Enhance Accessibility
In accessibility and HCI, wearable AR has been used to support people with diverse disabilities. For example, AR glasses can caption and visualize speech and sounds for deaf or hard of hearing (DHH) people (Findlater et al., 2019; Jain et al., 2018a, b, 2015; Miller et al., 2017; Olwal et al., 2020; Peng et al., 2018; Schipper and Brinkman, 2017), support hands-free interactions with screen displays for people with upper body motor impairments (McNaney et al., 2014; Malu and Findlater, 2014, 2015), offer speech support for people with aphasia (Williams et al., 2015), and provide social cue therapy for children with autism (Washington et al., 2017).
@@
Within the low-vision aid context, head-worn AR devices can selectively enhance users’ vision by interpreting their environment and tasks (Azenkot and Zhao, 2017; Zhao et al., 2017). For instance, prior research has developed AR systems that can capture real-time video feeds and apply image processing techniques (Massof and Rickman, 1992; Zhao et al., 2016; Deemer et al., 2018) to enhance visual information, such as edge enhancement (Massof et al., 1994; Kwon et al., 2012; Hwang and Peli, 2014), scene recoloring based on distance (Everingham et al., 1999; Hicks et al., 2013; van Rheede et al., 2015), and pixel remapping for visual field loss (Loshin and Juday, 1989; Gupta et al., 2018; Peli, 2001; Luo and Peli, 2006). However, while these solutions are beneficial for simple tasks like reading (Stearns et al., 2017; Huang et al., 2019; Deemer et al., 2018), they still lack semantic understanding of the scene and cannot effectively support more complex activities involving object interactions like cooking (Wang et al., 2023b).
@@
More recently, researchers have combined AR and CV to create scene-aware visual augmentations to assist LV people with more intricate activities like visual search (Zhao et al., 2016), stair navigation (Zhao et al., 2019), wayfinding (Zhao et al., 2020), obstacle maneuvering (Fox et al., 2023), button pressing (Lang and Machulla, 2021), and sports (Lee et al., 2023). Nonetheless, no prior systems have addressed the unique challenges of cooking involving dynamic tool interactions. Moreover, prior AR research for low vision primarily focuses on designing and evaluating visual augmentations. Therefore, they tend to oversimplify CV recognition in system development, such as by using QR codes (Zhao et al., 2016, 2019) or existing spatial mapping APIs (Zhao et al., 2020; Fox et al., 2023) to anchor augmentations to the real world. This approach neglects the technical challenges of building a real-time AI-powered AR system and the potential impacts of technological limitations (e.g., CV errors, system latency) on user experience.
Our research advances the field by introducing CookAR, a wearable AR system that recognizes and augments kitchen tool affordances in near real-time, enhancing safe and efficient interactions during cooking.
@@ 2.3. Affordance Segmentation
In contrast to most prior studies that augment objects as whole (Zhao et al., 2016; Fox et al., 2023; Lee et al., 2023), our research focuses on recognizing and augmenting tool affordances. Affordance is traditionally defined as “the opportunities for actions that objects offer, relative to the user’s ability to perceive and act on them” (Gibson, 1966, 1977, 2014). Highlighting object affordances can effectively guide human attention and actions (Garrido-Vásquez and Schubö, 2014; Tünnermann et al., 2015; Rehrig et al., 2022). Despite the prominence of affordance segmentation in robotics (Chu et al., 2019a, b; Nguyen et al., 2016, 2017; Bahl et al., 2023) and computer vision (Luddecke and Worgotter, 2017; Chuang et al., 2018; Luo et al., 2022), automatic affordance augmentation has received comparatively less attention and applications in the field of HCI. Notably, there is a lack of an egocentric affordance dataset specifically created for the needs of LV individuals. To address this gap, we first created a new dataset focused on kitchen tool affordances by selecting and annotating image frames from the egocentric Epic Kitchens video dataset (Damen et al., 2018) and fine-tuned an instance segmentation model. We then built an AR system that can segment and enhance tool affordance information.
@@ 3. System Implementation
To support safe and efficient LV hand-object interactions with kitchen tools, we designed and built CookAR, a wearable stereo AR prototype that recognizes and augments cooking tool affordances in near real-time. Unlike traditional enhancements that target objects as a whole, our prototype highlights their affordances (i.e., functional parts), facilitating identification and interactions with areas to grasp or avoid. To create a fully-functional wearable AR system, we needed to address both the computer vision problem of accurately recognizing object affordances in real-time and the HCI problem of designing and rendering suitable affordance augmentations. In this section, we describe our approach for each step, including (1) collecting and annotating a dataset focused on the affordances of kitchen tools; (2) fine-tuning an instance segmentation and recognition model on our dataset to detect these affordances; and (3) building a head-mounted AR system with a stereo camera to render visual augmentations on the recognized tool affordances. The labeled dataset and fine-tuned model weight are one contribution of our work and are open-sourced at: https://github.com/makeabilitylab/CookAR.